Last Fall Semester, 34% of freshmen who applied for football season tickets did not receive tickets for all home games. You randomly select a sample of 200 freshmen. What is the probability that more than 40% of those sampled did not receive tickets for all home games? Probability = (3 decimal places)
Answer #1

Using central limit theorem,

P( < p ) = P( Z < - p / sqrt ( p (1 -p ) / n)

so,

P( > 0.40) = P( Z > 0.40 - 0.34 / sqrt( 0.34 * 0.66 / 200) )

= P( Z > 1.7912)

= 1 - P( Z < 1.7912 )

= 0.963